ROTATIONAL SPECTRA AND AB INITIO POTENTIAL ENERGY SURFACE OF THE H_{2}-OCS VAN DER WAALS COMPLEX

\underline{KELLY HIGGINS}, Department of Chemistry, University of Minnesota, Minneapolis, Minnesota 55455; ZHENHONG YU, and WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, Massachusetts 02138.

The four-dimensional intermolecular potential energy surface for the H_{2}-OCS complex was obtained at the MP4 level. The potential gives a T-shaped global minimum with a distance of 3.2 Å between H_{2} and OCS, in which H_{2} is nearly parallel to OCS. Bound state calculations of paraH_{2}-OCS give predicted rotational constants of A = 22652, B = 5994, and C = 4611 MHz, in good agreement with the measured results from high-resolution infrared studies.^{a} The calculated binding energy of paraH_{2}-OCS is 75 cm^{−1}, almost four times greater than that of He-OCS.^{b} Preliminary bound state calculations of orthoH_{2}-OCS predict binding energies of 99 cm^{−1}for the Σ state and 64 cm^{−1}for the Π state.

a-type rotational transitions of two isotopomers of the orthoH_{2}-OCS complex were observed between 9 to 31 GHz. The spectral constants of orthoH_{2}-OC^{32}S are (B+C)/2 = 5113.372(23), (B-C)/2 = 580.337(40), and D_{J} = 2.118(2) MHz. The observed 1_{01}-0_{00} transition is very close to the expectation of paraH_{2}-OCS. This shows that the ground state of the complex is a Σ bound state with the T-shaped geometry. The H_{2} nuclear spin dipole-dipole coupling constant d_{HH}^{c} is 14.4(1) kHz which indicates a zero-point energy averaged angle of 45^{∘} of H_{2} with respect to the molecular a axis. Preliminary spectra of D_{2}-OCS show the existence of both orthoD_{2}-OCS and paraD_{2}-OCS complexes.

aJian Tang and A. R. W. McKellar, J. Chem. Phys. 116, 646 (2002).

bKelly Higgins and William Klemperer, J. Chem. Phys. 103, 1138 (1999).

cN. F. Ramsey, Molecular beam (Oxford University, London, 1953).